OpenStudy (anonymous):

An investor has $300,000 to invest in two types of investments. Type A pays 4% annually and type B pays 5% annually. To have a well-balanced portfolio, the investor imposes the following conditions. At least one-third of the total portfolio is to be allocated to type A investments and at least one-third of the portfolio is to be allocated to type B investments. What is the optimal amount that should be invested in each investment?

OpenStudy (anonymous):

0.04($300.000(1/3)=0.04(100.000) Plan A=$4000 0.05(300.000-100.000) 0.05(200.000) Plant B=$10.000

OpenStudy (anonymous):

Ok so I had something similar to this but the answers I can choose from don't match....

OpenStudy (anonymous):

you would want to have as much as possible in type B because it pays more. The most you could have is $200,000 due to the condition. So it is $100,000 to type A and $200,000 to type B

OpenStudy (anonymous):

My answer choices are: A.$100,000 in type A (4%), $200,000 in type B (6%) B. $0 in type A (4%), $300,000 in type B (6%) C.$200,000 in type A (4%), $100,000 in type B (6%) D.$300,000 in type A (4%), $0 in type B (6%) E.$110,000 in type A (4%), $190,000 in type B (6%)

OpenStudy (anonymous):

Yeah its A

OpenStudy (anonymous):

Ok so how did you come up with those numbers....

OpenStudy (anonymous):

because I have another problem: An investor has $600,000 to invest in two types of investments. Type A pays 7% annually and type B pays 9% annually. To have a well-balanced portfolio, the investor imposes the following conditions. At least one-third of the total portfolio is to be allocated to type A investments and at least one-third of the portfolio is to be allocated to type B investments. What is the optimal amount that should be invested in each investment? I am assuming then it is the same scenario so my choices would either be: a.$200,000 in type A (7%), $400,000 in type B (9%) b.$210,000 in type A (7%), $390,000 in type B (9%)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Nevermind I figured it out thanks anyways